# PLaID++: A Preference Aligned Language Model for Targeted Inorganic Materials Design

### v1

@@ Abstract
Discovering novel materials is critical for technological advancements such as solar cells, batteries, and carbon capture. However, the development of new materials is constrained by a slow and expensive trial-and-error process. To accelerate this pipeline, we introduce PLaID++, a Large Language Model (LLM) fine-tuned for stable and property-guided crystal generation. We fine-tune Qwen-2.5 7B to generate crystal structures using a novel Wyckoff-based text representation. We show that generation can be effectively guided with a reinforcement learning technique based on Direct Preference Optimization (DPO), with sampled structures categorized by their stability, novelty, and space group. By encoding symmetry constraints directly into text and guiding model outputs towards desirable chemical space, PLaID++ generates structures that are thermodynamically stable, unique, and novel at a $\sim$50% greater rate than prior methods and conditionally generates structures with desired space group properties. Our experiments highlight the effectiveness of iterative DPO, achieving $\sim$115% and $\sim$50% improvements in unconditional and space group conditioned generation, respectively, compared to fine-tuning alone. Our work demonstrates the potential of adapting post-training techniques from natural language processing to materials design, paving the way for targeted and efficient discovery of novel materials.
@@ 1 Introduction
Symmetry is a defining aspect of crystal structures. The set of rotations, reflections, inversions, and translations exhibited by a crystal lattice form its space group. A crystal’s space group and symmetries are not merely a mathematical construct but critical to many optical, electrical, and magnetic properties like piezoelectricity (Malgrange et al., 2014; Yang et al., 2005). One way to define a crystal’s structure from its symmetries is via Wyckoff positions, whereby one can specify a few key atomic coordinates and have the remaining atomic positions filled by applying symmetry operations (Hahn et al., 1983).
Other fields of machine learning have shown that general-purpose architectures that focus on scale and data expressivity outperform models with handcrafted, domain-specific constraints (Dosovitskiy, 2020; Kaplan et al., 2020). We contend that the same “Bitter Lesson” holds true for materials generation (Sutton, 2019). Rather than imposing explicit crystallographic constraints, we train an LLM to learn and exploit structural parameters from the data itself. By allowing the model to discover patterns implicitly through training on Wyckoff-based text representations, we see significant improvements in generation performance.
@@
By incorporating inherent crystal symmetries and the feedback of a Machine Learning Interatomic Potential (MLIP), we significantly increase the rate of stable, unique, and novel (S.U.N.) materials. Our experiments demonstrate that our method PLaID++ generates materials at state of the art stability rate and generates S.U.N materials at over a 20% higher rate than prior models while retaining the flexibility of natural language prompting. We also highlight the effectiveness of reinforcement learning for materials discovery, achieving $\sim$115% and $\sim$50% improvements in unconditional and space group conditioned generation, respectively, compared to fine-tuning alone. Our contributions are as follows:
@@
    We develop a novel, symmetry-informed text representation for crystal structures which is compact, performant, and physically-motivated.
@@
    We introduce an end-to-end framework for fine-tuning LLMs for material generation using Reinforcement-Learning from Interatomic Potentials (RLIP).
@@
    We demonstrate that fine-tuned LLMs outperform existing state-of-the-art models in generating novel and stable materials.
@@ 2 Related Work
On the other hand, symmetry-preserving non-LLM methods such as WyCryst (Zhu et al., 2024), Crystal-GFN (AI4Science et al., 2023), DiffCSP++ Jiao et al. (2023), and SymmCD Levy et al. (2024) enforce Wyckoff or space group constraints through templates or diffusion processes. While effective, these approaches either require predefined templates or incur high computational cost, and cannot flexibly target multiple objectives. In this paper, we explore using a compact Wyckoff text encoding with preference-based post-training, enabling both unconditional and conditional generation in a single unified pipeline.
@@ 3.1 Crystallography
Crystal structures are periodic arrangements of atoms defined by repeating unit cells. Each unit cell is fully described by lattice parameters—consisting of side lengths $(l_{1},l_{2},l_{3})$ and angles $(\theta_{1},\theta_{2},\theta_{3})$—and atomic positions within the cell, characterized by their fractional coordinates $(x_{i},y_{i},z_{i})$ and elemental identities $e_{i}$. Formally, a crystal structure $C$ can be expressed as a tuple:
$C=(l_{1},l_{2},l_{3},\theta_{1},\theta_{2},\theta_{3},e_{1},x_{1},y_{1},z_{1},\dots,e_{N},x_{N},y_{N},z_{N}).$ (1)
@@
In computational materials science, thermodynamic stability is a critical metric which can be described using the energy above hull ($E^{\text{hull}}$), which quantifies how energetically favorable a material is compared to known competing phases (Sun et al., 2016). The energy hull refers to a convex hull of formation energies in composition-energy space constructed by evaluating the formation energy of each computed phase and linking the lowest-energy points to form the lower convex envelope (Gruver et al., 2024). Hence, the $E^{\text{hull}}$ of a candidate material is defined as the distance between its computed formation energy per atom and the hull at the same composition.
Materials with $E^{\text{hull}}\leq 0$ eV/atom at a given temperature and pressure are considered thermodynamically stable, while those slightly above zero (typically $E^{\text{hull}}<0.1$ eV/atom) are considered metastable and potentially synthesizable (Jain et al., 2013). Compounds with lower $E^{\text{hull}}$ values have less thermodynamic driving force for decomposition, making them generally more likely to be synthesizable and thus have potential for real-world applications. Consequently, many generative models prioritize designing materials with low $E^{\mathrm{hull}}$ to maximize their physical viability.
Crystallography concerns not only the validity of atomic positions and crystal stability but also the uniqueness and novelty of generated structures. In practice, two crystals may have different symmetries yet could be considered chemically equivalent if their lattices or atomic sites overlap under tolerance thresholds. To address this, we rely on structure-based fingerprints such as Pymatgen’s StructureMatcher (Ong et al., 2013). StructureMatcher reduces candidate crystals to canonical forms (e.g., primitive and Niggli cells) and then checks atom-by-atom equivalence within length, angle, and site tolerances. Uniqueness and novelty therefore depend on the chosen reference set and tolerance hyperparameters. These metrics underpin the S.U.N. evaluation used throughout this work, complementing stability as a necessary criterion for materials discovery.
@@
Recent advances in reinforcement learning have enabled LLMs to align generated outputs more closely with human preferences (Bai et al., 2022). In general, these methods fine-tune a pre-trained LLM, $\pi_{ref}$ with feedback from a reward model $r_{\phi}(\cdot)$. The RL objective optimizes the model $\pi_{\theta}(y\mid x)$ to maximize the expected reward of generated responses $y\sim\pi_{\theta}(y\mid x)$, given prompt $x$. To preserve the diversity of response and prevent collapse to undesirable modes of the reward function, this is often regularized via an additional KL-divergence loss that prevents the model from straying too far from the initial model. This leads to the following general objective:
$\max_{\pi_{\theta}}\mathbb{E}_{x\sim D,y\sim\pi_{\theta}(y\mid x)}\left[r_{\phi}(x,y)\right]-\beta D_{\mathrm{KL}}\left[\pi_{\theta}(y\mid x)\|\pi_{\text{ref}}(y\mid x)\right].$ (2)
Here $D$ represents the distribution of prompts and the hyperparameter $\beta$ controls the strength of regularization. This objective can be optimized using REINFORCE or with reduced variance policy gradient such as proximal policy optimization (PPO) Schulman et al. (2017). This framework requires an explicit reward function, while actor-critic methods such as PPO additionally require a value model leading to substantially increased computational overhead.
In contrast, Direct Preference Optimization (DPO) (Rafailov et al., 2023) streamlines this process, using observed preference pairs of responses $(y_{w},y_{l})$, generated from $\pi_{ref}$, where $y_{w}$ denotes the preferred response. DPO optimizes an implicit reward function $r^{*}(x,y)$ by maximizing the likelihood of preferences under a Bradley-Terry preference model: $p(y_{w}\succ y_{l}\mid x)=\exp(r^{*}(y_{w}|x))/(\exp(r^{*}(y_{w}|x))+\exp(r^{*}(y_{l}|x)))$ . The authors show that a KL-regularized objective can be optimized without directly instantiating $r^{*}(x,y)$, leading to the following objective:
@@
Prior work has shown that DPO is effective in both RLHF (Rafailov et al., 2023) and scientific domains (Widatalla et al., 2024).
@@ 4.1 Base Model
The training objective is to maximize the likelihood that the model generates the correct token sequence representing stable crystal structures. We fine-tune two versions of the base model: one where the crystal structures are represented as text-based 3D coordinates and another based on a new text-based Wyckoff representation.
4.2 Wyckoff Representation
Crystal structures inherently follow symmetry constraints that can be represented by space groups and Wyckoff positions. To improve the model’s ability to generate valid and stable structures, we introduce a Wyckoff representation that encodes these symmetries directly into text. Unlike previous methods such as DiffCSP++ (Jiao et al., 2024) which rely on searching template structures for Wyckoff positions, our approach allows direct prediction of these sites within the generative process.
Because Wyckoff positions specify atomic site symmetries in a given space group, they significantly increase the possibility of feasible atomic arrangements. Our encoding method represents each crystal as a sequence of tokens as shown in Figure 2. The prompt used to train PLaID++ is the same for both unconditional and conditional generation except conditional generation includes the conditioning information on space group in blue. We specify the chemical formula, space group, lattice parameters, individual chemical elements, fractional coordinates, and Wyckoff site symmetries. Note that we include the chemical formula first to ensure self-consistency during generation.
@@ 4.3 Direct Preference Optimization (DPO)
Encoding symmetry improves structural validity; however, stability, uniqueness, and novelty remain critical for practical material discovery. To explicitly align our model toward generating crystals with desired properties, we employ an iterative form of Direct Preference Optimization (DPO) (Rafailov et al., 2023), a reinforcement learning technique that directly optimizes model outputs based on ranked preferences as highlighted by Equation 3. We jointly optimize our model on three explicit objectives: thermodynamic stability, novelty with respect to a reference crystal set, and space group conditioning.
@@
Uniqueness and Novelty   To incorporate novelty into our reward formulation, we construct preference pairs that distinguish between (stable & novel, stable & not novel). Incorporating uniqueness into a pairwise reward is non-trivial and presents many challenges. Namely, since uniqueness is defined at the group level (i.e. whether or not crystals within a set of rollouts are considered the “same” crystal as defined by Pymatgen’s StructureMatcher), pairwise reward computation of uniqueness is ambiguous. However, we can control sample diversity during rollouts via the sampling temperature. As temperature increases, uniqueness and novelty tend to increase, while stability and validity of generated crystals decrease. Based on this observation, we find that increasing the sampling temperature across successive iterations of DPO can prevent overoptimization for stability, preventing the model from collapsing on narrow regions of chemical space.
@@
5 Experiments
5.1 Setup
We trained our model on the well-established MP-20 dataset (Xie et al., 2021), a collection of 45,231 inorganic crystalline materials from the Materials Project (Jain et al., 2013). The dataset includes structures with up to 20 atoms, all of which are metastable. We follow the methodology of Gruver et al. (2024) by independently fine-tuning a pre-trained Qwen-2.5 7B model using 4-bit quantization and Low-Rank Adapters (LoRA) (Hu et al., 2021), implemented with PyTorch (Paszke et al., 2019) and Transformers (Wolf et al., 2020). Symmetry information for our Wyckoff representation is calculated by Pyxtal Fredericks et al. (2021). Following supervised fine-tuning, we apply DPO on the generated preference dataset to further guide generation towards stable structures. Full hyperparameters details are provided in Appendix 7.2.
To validate the stability and accuracy of our MLIP-based results, we performed Density Functional Theory (DFT) calculations on crystal samples generated by our flagship Qwen-based model fine-tuned with DPO and the Wyckoff encoding. Although accurate, DFT is computationally expensive, scaling typically as $O(N^{3})$ with respect to the number of atoms $N$. Therefore, in this work, in addition to performing DFT calculations, we also leverage machine learning interatomic potentials (MLIPs) such as eqV2 (Barroso-Luque et al., 2024) and eSEN (Fu et al., 2025) as efficient proxies for DFT. These MLIPs predict the relaxed atomic positions and energies with significantly reduced computational overhead. We specifically use eSEN for evaluation to ensure our generative models are guided by energetics closely aligned with DFT benchmarks, without incurring DFT’s computational expense during training. To verify our eSEN results, we additionally take a random 1,000 crystal subset of our generated crystals to compute stability using DFT. Detailed information regarding our DFT setup are in Appendix 7.2.
@@ 5.2 Metrics
Figure 3: Evolution of S.U.N. percentage of PLaID++ variants’ over DPO iterations. Reference lines represent S.U.N. rates from ADiT and FlowLLM’s flagship models.
@@ 5.3 Results
Unconditional Generation   We compare our model to seven prior methods, as reported from Joshi et al. (2025): CDVAE (Xie et al., 2021), DiffCSP (Jiao et al., 2023), FlowMM (Miller et al., 2024), FlowLLM (Sriram et al., 2024b), MatterGen-MP (Zeni et al., 2025), CrystalLLM (Gruver et al., 2024), and ADiT (Joshi et al., 2025). Our main results are presented in Table 1. On stability and S.U.N, the most important metrics, PLaID++ outperforms all prior methods. We also analyze the contribution of each training technique to our methodology. We see that our novel Wyckoff representation and iterative DPO both significantly improve S.U.N., resulting in our best PLaID++ model. For our best overall PLaID++ model, 22.27% of the relaxed structures are deemed stable, out of which 72% are novel, and 55% are unique, leading to a S.U.N. rate of 7.74%. PLaID++ achieves the highest stability rate and a $\sim$50% higher S.U.N rate than the best previous method.
@@
Conditional Generation   To test the ability of PLaID++ to generalize towards targeted structure synthesis, we measure the S.S.U.N. percentage for four model variants: (i) the 3D coordinate fine-tuned Qwen model (ii) the Wyckoff fine-tuned Qwen model, (iii) DPO conditioned only on space group pairs with no dynamic temperature or novelty reward,(iv) DPO trained jointly on space group, and stable and novel pairs, with dynamic temperature sampling (flagship PLaID++). As shown in Figure 4, applying DPO to space-group pairs increases S.S.U.N. by an average of 22% compared to the base Wyckoff model. Incorporating additional preference data from unconditional generation further boosts performance, resulting in a total increase of 47% over the base model—more than doubling the improvement from space-group conditioning data alone.
Groups with low training data occurrences like ($Amm2$, $Im2,P6_{3}/mmc$) show little improvement or sometimes slight degradation after DPO. This supports the elicitation hypothesis of reinforcement learning (Lambert, 2023), whereby RL amplifies existing valuable behaviors in the base model rather than teaching models new ones. With too few examples, neither the space-group nor the stability objective can generate a meaningful training signal for the model to learn from.
@@
Reinforcement Learning from Interatomic Potentials (RLIP)   Across multiple iterations of Direct Preference Optimization (DPO), we observe in Figure 3 that DPO successively improves the number of S.U.N. materials generated by Qwen-7B, demonstrating that reinforcement learning is an effective and robust method to improve generation performance. On unconditional and space group conditioned generation, we improve S.U.N. by $\sim$115% and $\sim$50% respectively compared to fine tuning alone. We believe this method could be scaled further given additional compute.
@@
From our ablation on temperature sampling in Figure 3, we highlight that removing dynamic temperature updates across DPO iterations leads to a loss of diversity, where the S.U.N. rate degrades even as stability continues to increase. This suggests that increasing temperature acts as an entropy regularizer which counteracts model collapse and encourages exploration.
@@
Joint Training Improves Performance   Table 1 and Figure 4 show that our DPO model trained together on the unconditional and space group conditioned generation tasks universally outperforms the DPO models trained individually on each respective task. This validates the efficacy of a universal foundational model that can be leveraged across multiple downstream objectives. The performance gains are particularly pronounced in conditional generation, where only a few hundred samples are available in the training set. This suggests that the additional training signal derived from unconditional generation feedback can be successfully transferred to conditional generation.
Space Group Distribution Analysis   We compare visualizations between the space group distribution of the MP-20 dataset we used to perform SFT on our base models, a sample of 10,000 crystals from Qwen SFT’d on the 3D coordinate representation, and a sample of 10,000 crystals we from our PLaID++ model trained on iterative DPO finetuning with our Wyckoff representation in Figure 5. We note that the flagship model favors generating out-of-train-distribution space groups such as $P2_{1}2_{1}2_{1}$ (19) and $Pnma$ (62), a consequence of our RLIP framework’s novelty incentive.
Space groups $Cm$ (8) and $P1$ (1) are prominent in the base 3D coordinate representation PLaID++ model’s crystals, but not in the flagship model. Space group Cm has a single symmetry operation (a glide plane), while space group $P1$ only has the identity operation. In addition, space groups $P6_{3}/mmc$ (194) and $I4mmm$ (139) are among the five most frequently generated space groups in the flagship model, even though they are not in the top five of the training set. This supports the hypothesis that the Wyckoff representation helps PLaID++ generate systems with complex symmetry requirements, skewing the distribution towards generating crystals with higher symmetries and ensuring a more even distribution across all space groups.
@@
Figure 5: A comparison between the three space group distributions allows us to extract valuable insights on how our PLaID variants affect model behavior.

### v2

@@ Abstract
Reinforcement Learning from Verifiable Rewards (RLVR) has emerged as a promising approach to improve correctness in LLMs, however, in many scientific problems, the objective is not necessarily to produce the correct answer, but instead to produce a diverse array of candidates which satisfy a set of constraints. We study this challenge in the context of materials generation. To this end, we introduce PLaID++ [^1] Our code is available at https://github.com/andaero/PLaID/tree/main, an LLM post-trained for stable and property-guided crystal generation. We find that performance hinges on our crystallographic representation and reward formulation. First, we introduce a compact, symmetry-informed Wyckoff text representation which improves computational efficiency and encourages generalization from physical priors. Second, we demonstrate that temperature scaling acts as an entropy regularizer which counteracts mode collapse and encourages exploration. By encoding symmetry constraints directly into text and guiding model outputs towards desirable chemical space, PLaID++ generates structures that are thermodynamically stable, unique, and novel at a $\sim$50% greater rate than prior methods and conditionally generates structures with desired space group properties. Our work demonstrates the potential of adapting post-training techniques from natural language processing to materials design, paving the way for targeted and efficient discovery of novel materials.
@@ 1 Introduction
Symmetry is a defining aspect of crystal structures. The set of rotations, reflections, inversions, and translations exhibited by a crystal lattice form its space group. A crystal’s space group and symmetries are not merely a mathematical construct but critical to many optical, electrical, and magnetic properties like piezoelectricity (Malgrange et al., 2014; Yang et al., 2005). One way to define a crystal’s structure from its symmetries is via Wyckoff positions, whereby one can specify a few key atomic coordinates and have the remaining atomic positions filled by applying symmetry operations (Hahn et al., 1983). We train an LLM to learn and exploit structural parameters from the symmetries in each crystal. By allowing the model to discover patterns implicitly through training on Wyckoff-based text representations, we see significant improvements in generation performance.
@@
By incorporating inherent crystal symmetries and the feedback of a Machine Learning Interatomic Potential (MLIP), we significantly increase the rate of stable, unique, and novel (S.U.N.) materials. Our experiments demonstrate that our method PLaID++ generates materials at a state of the art stability rate and generates S.U.N materials at a $\sim$50% higher rate than prior models while retaining the flexibility of natural language prompting. By using representations that encourage learning from physical priors and diversity-aware training, we also highlight the effectiveness of reinforcement learning for materials discovery, achieving $\sim$115% and $\sim$50% improvements in unconditional and space group conditioned generation, respectively, compared to fine-tuning alone. Our contributions are as follows:
@@
    We introduce Reinforcement-Learning from Interatomic Potentials (RLIP), a diversity-aware framework for fine-tuning LLMs for material generation.
@@
    Motivated by mode collapse after preference optimization on a coordinate-based representation, we develop a novel, symmetry-informed text representation for crystal structures which is compact, performant, and physically-motivated.
@@
    We demonstrate that unified training across conditional and unconditional generation tasks are mutually beneficial for data-sparse regimes. We achieve state-of-the-art results for generating novel and stable materials in both settings.
@@ 3.1 Crystallography
Crystal structures are periodic arrangements of atoms defined by repeating unit cells. Each unit cell is fully described by lattice parameters—consisting of side lengths $(l_{1},l_{2},l_{3})$ and angles $(\theta_{1},\theta_{2},\theta_{3})$—and atomic positions within the cell, characterized by their fractional coordinates $(x_{i},y_{i},z_{i})$ and elemental identities $e_{i}$.
@@
In computational materials science, thermodynamic stability is a critical metric which can be described using the energy above hull ($E^{\text{hull}}$), which quantifies how energetically favorable a material is compared to known competing phases (Sun et al., 2016). Materials with $E^{\text{hull}}\leq 0$ eV/atom at a given temperature and pressure are considered thermodynamically stable, while those slightly above zero (typically $E^{\text{hull}}<0.1$ eV/atom) are metastable and potentially synthesizable (Jain et al., 2013).
@@
RL methods align LLMs to human preferences by maximizing a learned reward while regularizing toward a reference model to prevent mode collapse (Bai et al., 2022). Actor–critic variants such as PPO reduce gradient variance but add overhead via explicit reward and value models (Schulman et al., 2017). Direct Preference Optimization (DPO) bypasses these components by fitting an implicit reward from preference pairs and optimizing a KL-regularized objective directly(Rafailov et al., 2023; Widatalla et al., 2024). DPO uses observed preference pairs of responses $(y_{w},y_{l})$, generated from $\pi_{ref}$, where $y_{w}$ denotes the preferred response. DPO optimizes an implicit reward function $r^{*}(x,y)$ by maximizing the likelihood of preferences under a Bradley-Terry preference model: $p(y_{w}\succ y_{l}\mid x)=\exp(r^{*}(y_{w}|x))/(\exp(r^{*}(y_{w}|x))+\exp(r^{*}(y_{l}|x)))$ . The authors show that a KL-regularized objective can be optimized without directly instantiating $r^{*}(x,y)$, leading to the following objective:
@@ 4.1 Base Model
The training objective is to maximize the likelihood that the model generates the correct token sequence representing stable crystal structures. We provide the prompt template used to fine-tune our model on both objectives in Figure 2. We fine-tune two versions of the base model: one where the crystal structures are represented as text-based 3D coordinates and another based on a new text-based Wyckoff representation.
@@ 4.2 Direct Preference Optimization (DPO)
To explicitly align our model toward generating crystals with desired properties, we employ an iterative form of Direct Preference Optimization (DPO) (Rafailov et al., 2023), a reinforcement learning technique that directly optimizes model outputs based on ranked preferences as highlighted by Equation 1. We jointly optimize our model on three explicit objectives: thermodynamic stability, novelty with respect to a reference crystal set, and space group conditioning.
@@
Uniqueness and Novelty   To incorporate novelty into our reward, we construct preference pairs that distinguish between (stable & novel, stable & not novel). Incorporating uniqueness into a pairwise reward is non-trivial and presents many challenges. Namely, since uniqueness is defined at the group level (i.e. whether or not crystals within a set of rollouts are considered the “same” crystal as defined by Pymatgen’s StructureMatcher (Ong et al., 2013)), pairwise reward computation of uniqueness is ambiguous. However, we can control sample diversity during rollouts via the sampling temperature. As temperature increases, uniqueness and novelty tend to increase, while stability and validity of generated crystals decrease. Based on this observation, we find that increasing the sampling temperature across successive iterations of DPO can prevent overoptimization for stability, preventing the model from collapsing on narrow regions of chemical space.
@@
4.3 Wyckoff Representation
When RLIP is naively trained on 3D coordinate representations, PLaID++ models tend to overfit preference pairs and undergo mode collapse, producing stable but repetitive crystals (See Section 5.3). We address this failure by introducing a novel Wyckoff representation.
Crystal structures inherently follow symmetry constraints that can be represented by space groups and Wyckoff positions. To improve the model’s ability to generate valid and stable structures, our Wyckoff representation encodes these symmetries directly into text. Unlike prior approaches such as DiffCSP++ (Jiao et al., 2024), which rely on searching template structures for Wyckoff sites, our method predicts them directly within the generative process.
Our representation is computationally efficient: on the MP-20 dataset, Wyckoff encodings average 185.5 tokens per crystal compared to 214.7 for coordinate-based strings, a 14% reduction that translates to better sample efficiency and faster training. Moreover, the format scales efficiently as our crystals grow more complex, since additional atoms at the same Wyckoff site require only a multiplicity update rather than new coordinate lines. By tying atoms to Wyckoff sites, a single change propagates through symmetry operations, forcing the model to generalize physical priors. Together, these properties make Wyckoff encoding a compact, symmetry-aware foundation that both improves efficiency and mitigates diversity collapse during RL fine-tuning.
@@ 4.3 Wyckoff Representation
5 Experiments
5.1 Setup
We trained our model on the well-established MP-20 dataset (Xie et al., 2021), a collection of 45,231 inorganic metastable crystalline materials from the Materials Project (Jain et al., 2013) with up to 20 atoms. We follow the methodology of Gruver et al. (2024) by independently fine-tuning a pre-trained Qwen-2.5 7B model using 4-bit quantization and Low-Rank Adapters (LoRA) (Hu et al., 2021), implemented with PyTorch (Paszke et al., 2019) and Transformers (Wolf et al., 2020). Symmetry information for our Wyckoff representation is calculated by Pyxtal Fredericks et al. (2021). Following supervised fine-tuning, we apply DPO on the generated preference dataset to further guide generation towards stable structures. Full hyperparameters details are provided in Appendix 7.2.
To validate the stability and accuracy of our results, we performed energy relaxations on samples generated by our flagship model fine-tuned with DPO and the Wyckoff encoding. Although accurate, Density Functional Theory (DFT) is computationally expensive, scaling typically as $O(N^{3})$ with respect to the number of atoms $N$. Therefore, we leverage machine learning interatomic potentials (MLIPs) such as eqV2 (Barroso-Luque et al., 2024) and eSEN (Fu et al., 2025) as efficient proxies for DFT. These MLIPs predict the relaxed atomic positions and energies with significantly reduced computational overhead. To verify our eSEN results, we compute stability using DFT on a random 1,000 crystal subset of our generated crystals. Detailed information regarding our DFT setup are in Appendix 7.2.
@@ 5.2 Metrics
Figure 3: Evolution of S.U.N. percentage of PLaID++ variants’ over DPO iterations. Reference lines represent S.U.N. rates from ADiT and FlowLLM’s flagship models. Left: Ablation over joint training. Right: Ablation over dynamic temperature.
@@ 5.3 Results
Unconditional Generation   We compare our model to seven prior methods, as reported from Joshi et al. (2025): CDVAE (Xie et al., 2021), DiffCSP (Jiao et al., 2023), FlowMM (Miller et al., 2024), FlowLLM (Sriram et al., 2024b), MatterGen-MP (Zeni et al., 2025), CrystalLLM (Gruver et al., 2024), and ADiT (Joshi et al., 2025). Our main results are presented in Table 1. On stability and S.U.N, the most important metrics, PLaID++ outperforms all prior methods. For our best overall PLaID++ model, 22.27% of the relaxed structures are deemed stable, out of which 72% are novel, and 55% are unique, leading to a S.U.N. rate of 7.74%. PLaID++ achieves the highest stability rate and a $\sim$50% higher S.U.N rate than the best previous method.
@@
Conditional Generation   To test the ability of PLaID++ to generalize towards targeted structure synthesis, we measure the S.S.U.N. percentage for four model variants: (i) the 3D coordinate fine-tuned Qwen model (ii) the Wyckoff fine-tuned Qwen model, (iii) DPO with data only from the conditional generation task (iv) DPO trained jointly on conditional and unconditional generation data (flagship PLaID++). As shown in Figure 4, applying DPO to space-group pairs increases S.S.U.N. by an average of 22% compared to the base Wyckoff model. Incorporating additional preference data from unconditional generation further boosts performance, resulting in a total increase of 47% over the base model—more than doubling the improvement from space-group conditioning data alone.
Groups with low training data occurrences (<< 400 samples each) like ($Amm2$, $Im2,P6_{3}/mmc$) show little improvement or sometimes slight degradation after DPO. This supports the elicitation hypothesis of reinforcement learning (Lambert, 2023), whereby RL amplifies existing valuable behaviors in the base model rather than teaching models new ones. With too few examples, neither the space-group nor the stability objective can generate a meaningful training signal for the model to learn from.
@@
Representation Choice is Critical to Prevent Mode Collapse
Metric It 1 It 2 It 3
S.U.N. (%) 2.93 2.23 1.69
Stability (%) 10.35 13.47 15.94
Table 2: PLaID 3D coordinate DPO’s S.U.N. and stability percentage across iterations of RLIP
In Table 2, we demonstrate the importance of our crystallographic representation. As we perform more iterations of DPO with the 3D coordinate representation, the diversity of the distribution quickly collapses and the model struggles to generalize from the DPO examples seen during training. Each iteration of the post-trained model has a progressively lower S.U.N. rate even as stability increases. This shows that with the incorrect representation, RL leads the model to memorize rather than generalize from reward feedback. We hypothesize that our Wyckoff representation works because of its higher information density, where local atom type or coordinate changes propagate to globally meaningful structure differences, expanding the diversity of valid structures the model can learn and generate. This is further reinforced by the ablation on the SFT base models in Table 1, where despite similar stability rates, the uniqueness and novelty of the Wyckoff representation is significantly higher, leading to a $\sim$25% increase in S.U.N.
@@
Reinforcement Learning from Interatomic Potentials (RLIP)   Across multiple iterations of Direct Preference Optimization (DPO), we observe in Figure 3 that DPO successively improves the number of generated S.U.N. materials, demonstrating that reinforcement learning is an effective and robust method to improve generation performance. On unconditional and space group conditioned generation, we improve S.U.N. by $\sim$115% and $\sim$50% respectively compared to fine tuning alone.
@@
From our ablation on temperature sampling in Figure 3, we highlight that removing dynamic temperature updates across DPO iterations leads to mode collapse, where the S.U.N. rate degrades even as stability continues to increase after 3-4 iterations. This suggests that increasing temperature acts as an entropy regularizer which encourages exploration and counteracts model collapse.
@@
Joint Training Improves Performance   Table 1 and Figure 4 show that our DPO model trained together on the unconditional and space group conditioned generation tasks outperforms the DPO models trained individually on each respective task. This validates the efficacy of a universal foundational model that can be leveraged across multiple downstream objectives. The performance gains are particularly pronounced in conditional generation, where only a few hundred samples are available in the training set. This suggests that the additional training signal derived from unconditional generation feedback can be successfully transferred to conditional generation.
PLaID++ Crystal Novelty Analysis   To better understand how reinforcement learning reshapes generative behavior, we visualize the compositional embedding space of S.U.N. crystals from the Wyckoff Base and PLaID++ models using UMAP (McInnes et al., 2018) (Figure 5). The left plot highlights a clear distributional shift: while both models occupy similar structural regions, PLaID++ expands into areas underrepresented in the base model, providing concrete evidence that RLIP enables local exploration around novel materials in our supervised baseline. In the right plot, we analyze the largest crystal cluster and observe that PLaID++ not only produces substantially more samples but also preferentially generates crystals containing P-block metalloids. Together, these visualizations demonstrate that PLaID++ is not merely memorizing preferences but uncovering out-of-distribution materials with meaningful structural diversity.
@@
Figure 5: Left: UMAP visualizations of compositional embeddings of S.U.N. crystals generated by our PLaID++ and 3D Coord Base models. Right: Identical data, with crystals containing P-block metalloids being highlighted in red and those without P-block metalloids being in blue.